### v2

Abstract: This paper investigates the structure of $K_{2}$-groups for certain quotient rings of the integral group ring $\mathbb{Z}[G]$. Let $G$ be a finite abelian $p$-group with $p$-rank $r$, let $\Gamma$ be the maximal $\mathbb{Z}$-order of $\mathbb{Q}[G]$, and let $\widetilde{G}$ denote the sum of all elements of $G$ in the group ring. By employing the framework of Kähler differentials, we first determine that the relative $K$-group $)K_{2}(\mathbb{F}_{p}[G],(\widetilde{G}$)) is an elementary abelian $p$-group of rank $r$ when $|G|>2$. Building upon this result, we establish an explicit isomorphism for $r>1$:
$).K_{2}(\mathbb{Z}[G]/(|G|\Gamma\cap p\mathbb{Z}[G]))\cong K_{2}(\mathbb{Z}[G]/|G|\Gamma)\oplus K_{2}(\mathbb{F}_{p}[G],(\widetilde{G}$)).
Key words: $K_{2}$-groups; group rings; Dennis-Stein symbols; Kähler differentials
[…]
This paper is a continuation of our previous work [1]. Let $G$ be a finite abelian $p$-group with $p$-rank $r$, and let $\Gamma$ denote the maximal $\mathbb{Z}$-order in the group algebra $\mathbb{Q}[G]$. Set $J=|G|\Gamma$. By Proposition 2.2 in [2], $\Gamma$ is isomorphic to a direct product of rings of algebraic integers, with the inclusions $J\subseteq\mathbb{Z}[G]\subseteq\Gamma$ holding.
In [3], the group $K_{2}(\mathbb{Z}[G]/J)$ was investigated to establish a lower bound for the order of $K_{2}(\mathbb{Z}[G])$. It was shown that $K_{2}(\mathbb{Z}[G]/J)$ is trivial if and only if the order of $G$ is square-free. However, determining the explicit structure of $K_{2}(\mathbb{Z}[G]/J)$ remains an open problem, largely due to the complexity of $\Gamma$ (see [3, Lemma 3.3]). In our previous work [1], we provided computational evidence and nontrivial examples for groups $G$ of small order.
Let $I=J\cap p\mathbb{Z}[G]$ and $\widetilde{G}=\sum_{g\in G}g$. These definitions give rise to the following Cartesian square:
$\mathbb{Z}[G]/I$$\mathbb{Z}[G]/J$$\mathbb{F}_{p}[G]$${\mathbb{F}_{p}[G]/(\widetilde{G})}$ (0.1)
In the present paper, we establish an explicit relation between the $K_{2}$-groups of the rings lying in (0.1). The rest of the article is organized as follows: Section 1 is devoted to the necessary preliminaries.
In Section 2, we employ the framework of Kähler differentials to prove two key results for $|G|>2$, which are then applied to determine the explicit structure of $)K_{2}(\mathbb{F}_{p}[G],(\widetilde{G}$)) (see Lemma 3.1). Furthermore, Theorem 2.3 is used to establish the excision property for $K_{2}$ with respect to the cartesian square (0.1), as described in Proposition 3.3.
As a consequence, in Section 3, we show that the following exact sequences of finite abelian $p$-groups split when $r>1$:
$1\rightarrow K_{2}(\mathbb{Z}[G]/I,J/I)\rightarrow K_{2}(\mathbb{Z}[G]/I)\rightarrow K_{2}(\mathbb{Z}[G]/J)\rightarrow 1,$
$1\rightarrow K_{2}(\mathbb{F}_{p}[G],(\widetilde{G}))\rightarrow K_{2}(\mathbb{F}_{p}[G])\rightarrow K_{2}(\mathbb{F}_{p}[G]/(\widetilde{G}))\rightarrow 1.$
Combining this with the aforementioned excision isomorphism, we establish the decomposition (see Theorem 3.5):
[…]
This result, obtained by extending the methods of [3] to a more general setting, provides a structural decomposition for the $K_{2}$-groups of the quotient rings of $\mathbb{Z}[G]$ in square (0.1), and extends our previous computational studies to a general theoretical framework.
[…]
According to [4, Chapter III, §5], if $I$ is a radical ideal of a commutative ring $R$, then 1ab1-ab is a unit for all $a\in R,b\in I$, and the relative $K_{2}$-group $K_{2}(R,I)$ is generated by the Dennis–Stein symbols $\langle a,b\rangle$ for $(a,b)\in(R\times I)\cup(I\times R)$. These symbols satisfy the following defining relations (written multiplicatively):
  1. (DS1)
    $\langle a,b\rangle\langle b,a\rangle=1$
  2. (DS2)
    $\langle a,b\rangle\langle a,c\rangle=\langle a,b+c-abc\rangle$
  3. (DS3)
    $\langle a,bc\rangle=\langle ab,c\rangle\langle ac,b\rangle$  (where at least one of $a,b,c$ belongs to $I$)
Remark 1.1.
In the present study, the ideal $I$ satisfies $I^{2}=0$, which naturally ensures the radical condition. Under this setting, the symbol $\langle a,b\rangle$ used here (following [4]) coincides with the notation $\langle-a,b\rangle^{-1}$ in earlier literature such as [5]. Furthermore, the term abcabc in (DS2) vanishes, which reduces the relation to $\langle a,b+c\rangle=\langle a,b\rangle\langle a,c\rangle$. By the skew-symmetry from (DS1), the symbol $\langle a,b\rangle$ is additive in both components. This linearity simplifies the structure of $K_{2}(R,I)$ to a framework compatible with Kähler differentials, as seen in Theorem 2.3.
The following lemma is established by iterated applications of (DS1) and (DS3).
Lemma 1.2.
[6, p. 255] Let $\alpha_{0},\dots,\alpha_{l}$ be elements of a ring such that $1-\alpha_{0}\cdots\alpha_{l}$ is invertible. Let $\hat{\alpha}_{i}=\alpha_{0}\cdots\alpha_{i-1}\alpha_{i+1}\cdots\alpha_{l}$. Then
$1=\langle 1,\alpha_{0}\cdots\alpha_{l}\rangle=\prod_{i=0}^{l}\langle\alpha_{i},\hat{\alpha}_{i}\rangle.$
Given a ring homomorphism $\phi:R\to S$ that endows $S$ with an $R$-algebra structure, the module of Kähler differentials of $S$ over $R$ which we denoted by $\Omega_{S/R}$, is the $S$-module generated by the set $\{\mathrm{d}(f)|f\in S\}$ subjecting to the following relations:
  1. (i)
    $\mathrm{d}(s_{1}s_{2})=s_{1}\mathrm{d}(s_{2})+s_{2}\mathrm{d}(s_{1})$;
  2. (ii)
    $\mathrm{d}(r_{1}s_{1}+r_{2}s_{2})=s_{1}\mathrm{d}(s_{2})+s_{2}\mathrm{d}(s_{1})$;
  3. (iii)
    $1\cdot\mathrm{d}(r)=0$ for all $r\in R$.
We will abbreviate $\Omega_{S/\mathbb{Z}}$ as $\Omega_{S}$ if there is no confusion. The following exact sequence is known as the conormal sequence.
Lemma 1.3.
[7, Proposition 16.12] If $\pi:S\rightarrow T$ is an epimorphism of $R$-algebras, with kernel $I$, then there is an exact sequence of $T$-modules
$I/I^{2}\xrightarrow{\mathrm{d}}T\otimes_{S/R}\Omega_{S/R}\xrightarrow{\mathrm{D}\pi}\Omega_{T/R}\rightarrow 0,$
where the right-hand map is given by $\mathrm{D}\pi:c\otimes\mathrm{d}b\mapsto c\mathrm{d}b$ and the left-hand map takes the class of $f$ to $1\otimes\mathrm{d}f$.
[…]
In this section, we provide two technical results necessary for the following sections. Lemma 2.1 is a special case of [8, Lemma 1], for which we offer an alternative proof. This result will be applied to establish Lemma 3.1. Additionally, Theorem 2.3 is required to prove Proposition 3.3.
Lemma 2.1.
Let $G=\langle g_{1}\rangle\times\cdots\times\langle g_{r}\rangle$ be a finite abelian $p$-group. Then $\Omega_{\mathbb{F}_{p}[G]}$ is a free $\mathbb{F}_{p}[G]$-module with basis $\mathrm{d}g_{1},\cdots,\mathrm{d}g_{r}$.
Proof.
Let $p^{n_{i}}$ be the order of $g_{i},1\leq i\leq r$. Obviously, the $\mathbb{F}_{p}$-algebra $S=\mathbb{F}_{p}[x_{1},\cdots,x_{r}]$ maps onto $T=\mathbb{F}_{p}[G]$, and the kernel $I$ is generated by $x_{1}^{p^{n_{1}}}-1,\cdots,x_{r}^{p^{n_{r}}}-1$. Since $\mathrm{d}(I)=0$, by Lemma 1.3, we have
$\Omega_{\mathbb{F}_{p}[G]}\cong\mathbb{F}_{p}[G]\otimes_{\mathbb{F}_{p}[x_{1},\cdots,x_{r}]}\Omega_{\mathbb{F}_{p}[x_{1},\cdots,x_{r}]},$
then our lemma follows from the fact that $\Omega_{\mathbb{F}_{p}[x_{1},\cdots,x_{r}]}$ is a free $\mathbb{F}_{p}[x_{1},\cdots,x_{r}]$-module with basis $\mathrm{d}x_{1},\cdots,\mathrm{d}x_{r}$.
Theorem 2.2.
Let $A$ be a commutative ring, $I$ an ideal of $A$. Suppose $J=(b_{1}\cdots b_{r})A+I$ is an ideal of $A$ with $r>1$ such that $b_{i}J\subset I$ for each $i$, then
$K_{2}(A/I,J/I)\cong J/I\otimes_{A/I}\Omega_{A/I}.$
Proof.
Let $p^{n_{i}}$ be the order of $g_{i}$ for $1\leq i\leq r$. Consider the surjective $\mathbb{F}_{p}$-algebra homomorphism from $S=\mathbb{F}_{p}[x_{1},\cdots,x_{r}]$ to $T=\mathbb{F}_{p}[G]$. The kernel $I$ is generated by $x_{1}^{p^{n_{1}}}-1,\cdots,x_{r}^{p^{n_{r}}}-1$. Since $\mathrm{d}(I)=0$, by Lemma 1.3, we have
$\Omega_{\mathbb{F}_{p}[G]}\cong\mathbb{F}_{p}[G]\otimes_{\mathbb{F}_{p}[x_{1},\cdots,x_{r}]}\Omega_{\mathbb{F}_{p}[x_{1},\cdots,x_{r}]}.$
The result follows from the fact that $\Omega_{\mathbb{F}_{p}[x_{1},\cdots,x_{r}]}$ is a free $\mathbb{F}_{p}[x_{1},\cdots,x_{r}]$-module with basis $\mathrm{d}x_{1},\cdots,\mathrm{d}x_{r}$.
Theorem 2.3.
Let $A$ be a commutative ring and $I$ be an ideal of $A$. Suppose $J=(b_{1}\cdots b_{r})A+I$ is an ideal of $A$ with $r>1$ such that $b_{i}J\subset I$ for each $i$. Then
$K_{2}(A/I,J/I)\cong J/I\otimes_{A/I}\Omega_{A/I}.$
Proof.
Since $b_{i}J\subset I$ for each $i$, it follows that $J^{2}\subset I$; thus, for any $(J/I)^{2}=0$. According to Theorem 1.3 in [5], there exists an exact sequence of $A/J$-modules:
$J/I\otimes_{A/J}J/I\xrightarrow{\psi}K_{2}(A/I,J/I)\xrightarrow{\rho}J/I\otimes_{A/J}\Omega_{A/J}\rightarrow 0,$
where $K_{2}(A/I,J/I)$ is generated by the Dennis-Stein symbols $\langle a,b\rangle$ with $a$ or $b$ in $J/I$. The map $\psi$ is defined by $a\otimes b\mapsto\langle a,b\rangle$, and by Remark 1.1 it is straightforward to verify that $\psi$ is a bilinear map. The map $\rho$ is given by
$\rho\langle a,b\rangle=\begin{cases}a\otimes\mathrm{d}b,&\text{if }a\in J/I;\\ -b\otimes\mathrm{d}a,&\text{if }b\in J/I.\end{cases}$
For $s,t\in A$, let $\alpha_{0}=s$ and $\alpha_{i}=b_{i}$ for $1\leq i\leq r$. We have
$\quad\psi(b_{1}b_{2}\cdots b_{r}s\otimes b_{1}b_{2}\cdots b_{r}t)$
$=\langle b_{1}b_{2}\cdots b_{r}s,b_{1}b_{2}\cdots b_{r}t\rangle$
$=\langle s,(b_{1}b_{2}\cdots b_{r})^{2}t\rangle\prod_{i=1}^{r}\langle b_{i},(b_{1}b_{2}\cdots b_{r})^{2}st/b_{i}\rangle$ (by Lemma 1.2)
$=1.$ (since $b_{i}J\subset I$)
The second-last equality follows by regarding the first entry of $\langle b_{1}\cdots b_{r}s,b_{1}\cdots b_{r}t\rangle$ as a product of $r+1$ factors and leaving the second entry unchanged. The last equality holds because $r>1$ and $b_{i}J\subset I$ for each $i$, so that the elements $(b_{1}\cdots b_{r})^{2}t$ and $(b_{1}\cdots b_{r})^{2}st/b_{i}$ all lie in $I$. In the relative group $K_{2}(A/I,J/I)$, these elements represent 0 in $A/I$, which implies that every Dennis-Stein symbol in the product evaluates to the identity element $1$.
Thus, $\operatorname{Im}(\psi)=1$, it follows that $K_{2}(A/I,J/I)\cong J/I\otimes_{A/J}\Omega_{A/J}$. To complete the proof, it suffices to show that $J/I\otimes_{A/J}\Omega_{A/J}\cong J/I\otimes_{A/I}\Omega_{A/I}$. According to Lemma 1.3, there is an exact sequence of $A/J$-modules:
$J/I\xrightarrow{\delta}A/J\otimes_{A/I}\Omega_{A/I}\rightarrow\Omega_{A/J}\rightarrow 0.$
Tensoring the sequence above with $J/I$, we obtain
$J/I\otimes_{A/J}J/I\xrightarrow{\delta^{*}}\Omega_{A/I}\otimes_{A/I}J/I\rightarrow J/I\otimes_{A/J}\Omega_{A/J}\rightarrow 0,$
where $\delta^{*}$ is a bilinear map. For $s,t\in A$, we have
$\delta^{*}(b_{1}b_{2}\cdots b_{r}s\otimes b_{1}b_{2}\cdots b_{r}t)$ $=b_{1}b_{2}\cdots b_{r}t\otimes\mathrm{d}(b_{1}b_{2}\cdots b_{r}s)$
$=(b_{1}b_{2}\cdots b_{r})^{2}t\otimes\mathrm{d}s+\sum_{i=1}^{r}(b_{1}b_{2}\cdots b_{r})^{2}t/b_{i}\otimes\mathrm{d}b_{i}$
$=0\otimes\mathrm{d}s+\sum_{i=1}^{r}0\otimes\mathrm{d}b_{i}$
$=0.$
Hence, $J/I\otimes_{A/J}\Omega_{A/J}\cong J/I\otimes_{A/I}\Omega_{A/I}$. This completes the proof of the theorem.
[…]
In this section, we extend several results from [3, Section 4] concerning $K_{2}$-relations for rings lying in the Cartesian square (0.1).
Suppose $G=C_{p^{n_{1}}}\times\cdots\times C_{p^{n_{r}}}$ is a finite abelian $p$-group with generators $g_{1},\dots,g_{r}$, and let $\widetilde{G}=\sum_{g\in G}g$ denote the sum of all group elements. Let $\Gamma$ be the maximal $\mathbb{Z}$-order in the rational group algebra $\mathbb{Q}[G]$. We define $J=|G|\Gamma$ and $I=J\cap p\mathbb{Z}[G]$ as ideals of $\mathbb{Z}[G]$. As the structure of $\Gamma$ is well understood [3, Lemma 3.3], we have the relations $J=I+\widetilde{G}\mathbb{Z}$ and $p\widetilde{G}\in I$ hold.
Lemma 3.1.
Let $G$ be as above with $|G|>2$, and set $x_{i}=g_{i}-1$. Then $)K_{2}(\mathbb{F}_{p}[G],(\widetilde{G}$)) is an elementary abelian $p$-group of rank $r$ with basis $\{\langle x_{i},\prod_{j=1}^{r}x_{j}^{p^{n_{j}}-1}\rangle\mid 1\leq i\leq r\}.$
Proof.
Note that $x_{i}G=0$ for all $1\leq i\leq r$, and the following identity holds in $\mathbb{F}_{p}[G]$:
$\widetilde{G}=\prod_{j=1}^{r}\left(\sum_{i=0}^{p^{n_{j}}-1}(g_{j})^{i}\right)=\prod_{j=1}^{r}(g_{j}-1)^{p^{n_{j}}-1}=\prod_{j=1}^{r}x_{j}^{p^{n_{j}}-1}.$
Thus, $\widetilde{G}$ has degree greater than 1 if and only if $|G|>2$. Since $|G|>2$ holds, according to Theorem 2.3, there exists an isomorphism
$K_{2}(\mathbb{F}_{p}[G],(\widetilde{G}))\cong(\widetilde{G})\otimes_{\mathbb{F}_{p}[G]}\Omega_{\mathbb{F}_{p}[G]}.$
By Lemma 2.1, $\Omega_{\mathbb{F}_{p}[G]}$ is a free $\mathbb{F}_{p}[G]$-module of rank $r$. Since $p\widetilde{G}=0$ in $\mathbb{F}_{p}[G]$, the group $)K_{2}(\mathbb{F}_{p}[G],(\widetilde{G}$)) has exponent $p$, which implies that it is an elementary abelian $p$-group of rank $r$. Consequently, by Lemma 2.1, the set $\{(\prod_{j=1}^{r}x_{j}^{p^{n_{j}}-1})\otimes\mathrm{d}x_{i}\mid 1\leq i\leq r\}$ constitutes a basis for $(\widetilde{G})\otimes_{\mathbb{F}_{p}[G]}\Omega_{\mathbb{F}_{p}[G]}$. Via the map $\rho$ defined in Theorem 2.3, this basis corresponds to a basis of $)K_{2}(\mathbb{F}_{p}[G],(\widetilde{G}$)), and the result follows.
Remark 3.2.
$)K_{2}(\mathbb{F}_{p}[G],(\widetilde{G}$)) is trivial when $G=C_{2}$, as noted in [3, Lemma 4.1].
Proposition 3.3.
Let $G,I,J$ be as above with $|G|>2$. Then
$).K_{2}(\mathbb{Z}G/I,(\widetilde{G}))\cong K_{2}(\mathbb{F}_{p}[G],(\widetilde{G}$)).
Proof.
Let $A=\mathbb{Z}[G]$, $I_{1}=I$, $I_{2}=pA$, $J_{1}=J$, and $J_{2}=(\widetilde{G})A+I_{2}$. For each $1\leq j\leq r$, define $b_{j}=\sum_{i=1}^{p^{n_{j}}}g_{j}^{i}$. Then we have $J_{1}/I_{1}=\widetilde{G}(A/I_{1})$ and $J_{2}/I_{2}=\widetilde{G}(A/I_{2})$. Since $\widetilde{G}=\prod_{j=1}^{r}b_{j}$ and $p(J_{1}/I_{1})=0$, it follows from Theorem 2.3 that $K_{2}(A/I_{1},J_{1}/I_{1})$ is an elementary abelian $p$-group of rank at most $r$, and that there exists a surjective map from $K_{2}(A/I_{1},J_{1}/I_{1})$ to $K_{2}(A/I_{2},J_{2}/I_{2})$. By Lemma 3.1, the $p$-rank of $K_{2}(A/I_{2},J_{2}/I_{2})$ is exactly $r$. Comparing these ranks, we conclude that $K_{2}(A/I_{1},J_{1}/I_{1})$ must also be of rank $r$, and therefore the aforementioned surjection is an isomorphism.
Remark 3.4.
Since $I_{1}=J_{1}\cap I_{2}$, we have $(J_{1}/I_{1})\cap(I_{2}/I_{1})=0$. Using this fact, by Theorem 14 in [9], we obtain an exact sequence:
$(J_{1}/I_{1})/(J_{1}/I_{1})^{2}\otimes_{A^{e}}(I_{2}/I_{1})/(I_{2}/I_{1})^{2}\xrightarrow{\psi}K_{2}(A/I_{1},J_{1}/I_{1})\rightarrow K_{2}(A/I_{1},J_{2}/I_{2})\rightarrow 1,$
where $A^{e}=A\otimes_{\mathbb{Z}}A^{op}$ and the map $\psi$ is defined by $\psi(\overline{a}\otimes_{A^{e}}\overline{b})=\langle a,b\rangle$. In our case, since $(J_{1}/I_{1})^{2}=0$, Remark 1.1 ensures that the Dennis-Stein symbol $\langle a,b\rangle$ is additive in both components. Given that $p\widetilde{G}\in(J_{1}/I_{1})\cap(I_{2}/I_{1})=0$ in $A/I_{1}$, for any $s,t\in A$, we have:
$\psi(\overline{\widetilde{G}s}\otimes_{A^{e}}\overline{pt})=\langle\widetilde{G}s,pt\rangle=\langle p\widetilde{G}s,t\rangle=\langle 0,t\rangle=1.$
Hence, the image of $\psi$ is trivial. This yields an alternative proof of the isomorphism.
Theorem 3.5.
Let $G,I,J$ be as above with $r>1$. Then
$).K_{2}(\mathbb{Z}[G]/I)\cong K_{2}(\mathbb{Z}[G]/J)\oplus K_{2}(\mathbb{F}_{p}[G],(\widetilde{G}$)).
Proof.
The Cartesian square (0.1) gives rise to a natural commutative diagram with exact rows:
[…]
where the surjectivity of the maps follows from [2, Proposition 1.1]. The homomorphism $i$ is injective: each generator of $)K_{2}(\mathbb{F}_{p}[G],(\widetilde{G}$)) has order $p$, and each corresponding image of $i$ denoted by the same literal, also has order $p$ and lies in a basis for $K_{2}\left(\mathbb{F}_{p}[G]\right)$ when $r>1$ (see [10]). As a result, the bottom exact sequence splits: the monomorphism $i$ has a section, which we denote by $j$. Since $g_{1}$ is an isomorphism by Proposition 3.3, the commutativity of the first square implies that $f_{1}$ is injective. Consequently, $g_{1}^{-1}jg_{2}$ is a section of $f_{1}$, showing that $f_{1}$ is a split injection. The top exact sequence therefore splits, and we obtain
$).K_{2}(\mathbb{Z}[G]/I)\cong K_{2}(\mathbb{Z}[G]/J)\oplus K_{2}(\mathbb{F}_{p}[G],(\widetilde{G}$)).
Remark 3.6.
By the Snake Lemma, we have $\mathrm{Ker}(g_{2})\cong\mathrm{Ker}(g_{3})$. In addition, the corresponding square for $K_{2}$-groups in (0.1) is Cartesian in the category of groups.
[…]
We thank the referees for their time and comments.
[…]
  • [1] Zhang, Y. and Tang, G.P., $K_{2}$ for a kind of special finite group rings, J. Univ. Chinese Acad. Sci., 2018, 35(6):721.
  • [2] Alperin, R.C., Dennis, R.K. and Stein, M.R., $SK_{1}$} of finite abelian groups. I, Invent. Math., 1985, 82(1):1–18.
  • [3] Chen, H., Gao, Y.B. and Tang, G.P., $K_{2}$} of a Quotient Ring of $\mathbb{Z}{G}$, Comm. Algebra, 2014, 42(4):1571–1581.
  • [4] Weibel, C.A., The $K$-book: An Introduction to Algebraic $K$-theory, volume 145 of Grad. Stud. Math.. Amer. Math. Soc., Providence, RI, 2013.
  • [5] Weibel, C.A., $K_{2}$}, $K_{3}$} and nilpotent ideals, J. Pure Appl. Algebra, 1980, 18(3):333–345.
  • [6] Alperin, R.C., Dennis, R.K., Oliver, R. and Stein, M.R., $SK_{1}$} of finite abelian groups. II, Invent. Math., 1987, 87(2):253–302.
  • [7] Eisenbud, D., Commutative Algebra: With a View Toward Algebraic Geometry. Springer, 2013.
  • [8] Magurn, B.A., Explicit $K_{2}$} of some finite group rings, J. Pure Appl. Algebra, 2007, 209(3):801–811.
  • [9] Keune, F., The relativization of $K_{2}$}, J. Algebra, 1978, 54(1):159–177.
  • [10] Zhang, H., Tang, G.P. and Liu, H., Some remarks on $K_{2}$} and $K_{3}$} of finite abelian group algebras, J. Algebra Appl., 2019, 18(5):1950094.